OpenStudy (anonymous):

even or odd? y=x^3 - x

OpenStudy (anonymous):

Even

OpenStudy (anonymous):

y = x(x^2-1) y = x(x+1)(x-1) so if x is odd, then odd(even)(even) = even. if x is even, even(odd)even = even. @Luis_Rivera its (-1)^3 - (-1) = -1 + 1 = 0 which is not treated as odd or even but u plug in any other values u get even number

OpenStudy (anonymous):

y=x(x^2-1)=x(x+1)(x-1) hence x can be 0,-1 or +1 and hence can be odd or even :D

OpenStudy (anonymous):

sry it was x even then even(odd)(odd) which is actually even,

OpenStudy (anonymous):

yep ^

OpenStudy (anonymous):

@sriramkumar odd x even x odd = even. 3 x 2 x 3 = 18 = even. OK? Y will hence be even for any value of x.

OpenStudy (anonymous):

provided u take x > 0, any odd negative value for x and u get a negative even number actually

OpenStudy (anonymous):

ohh... chill dude... you've hit the right target :D @luckythebest